OpenStudy (anonymous):

Solve the equation by factoring. 12x^2+12x=24 x= (smaller value) x= (larger value)

OpenStudy (anonymous):

subtract 24 12x^2+12x-24=0 factor out 12 12(x^2+x-2)=0 factor the (x^2+x-2)=0 get 12(x+1)(x-2)=0 plug the +1 and the -2 back into the original equation. success :)